# MARS: A neurosymbolic approach for interpretable drug discovery

### v2

@@ 3.3 Weight updates via 2-hop joint probability
Here, the probability of each metapath-based rule is computed as the joint probability of its fragments. For example, the $P_{2H}$ metric for the above metapath-based rule would be computed as $P_{2H}(M_{example})=p(Fragment1)\times p(Fragment2)\times p(Fragment3)$. The probability of each fragment is determined by the ratio of its observed and expected frequencies (see Appendix A.3). We note that, while the fragments do not necessarily represent independent events, the way in which $P_{2H}$ is computed imposes this assumption. To compute the joint probability for dependent events, we require the conditional probabilities of the dependent events (Russell & Norvig, 2010). This can, in theory, be computed, but it becomes more complex with metapath length. For simplicity’s sake, we assume independence, and the $P_{2H}$ metric serves as an approximation for the empirical probabilities of metapath-based rules. Further formalization and pseudocode are in Appendix A.3. Ultimately, MARS with $P_{2H}$ updates ($\text{MARS}_{P_{2H}}$) uses all information from successful trajectories.
3.4 Datasets: MoA-net and its variants
We design our KG, MoA-net [^2] https://github.com/laurendelong21/MoA-Net, specifically for MoA prediction. MoA-net consists of drugs, proteins, and BPs (Fig. 1-A). We assemble it using the causal relations between drugs and proteins from Custom KG (Rivas-Barragan et al., 2020) and OpenBioLink KG (Breit et al., 2020). The BP nodes come from molecular function annotations in UniProt (Consortium, 2015).
To predict drug-BP triples, which are unique to MoA-net, we make use of publicly available functional and biochemical assays in ChEMBL (v33), an open access database of bioactive compounds (Gaulton et al., 2012). Of the 1,622 drug-BP triples obtained, 48 also had known MoAs in DrugMechDB (Gonzalez-Cavazos et al., 2023), a manually-curated compendium of MoAs. Between the three node types, we define five unique edge types, or relations, shown in Fig. 1-B. We also include all inverse relations, running in the opposite direction of causality.
Using the hetnetpy package (Himmelstein et al., 2021), we extract all metapaths (see Section 3.1) from MoA-net which we considered to be valid MoAs: those comprising directed, mechanistic paths between drug and BP nodes (see Appendix A.1). Although other metapaths are possible, we exclude them from our set of metapath-based rules. Based on MoAs found in DrugMechDB, we limit metapaths to a maximum length of four relations (or hops).
Figure 1: A) MoA-net schema. B) Distribution of node and relationship types of MoA-net. Note that the $y$-axis has a logarithmic scale.
3.5 Evaluation
We conduct an extensive benchmark of our method against nine different baseline KG embedding (KGE) models, two state-of-the-art NeSy methods, and one network measure. Based on previous biomedical KG benchmarks (Rivas-Barragan et al., 2022), the KGE models include RESCAL (Nickel et al., 2011), TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ER-MLP (Dong et al., 2014), ComplEx (Trouillon et al., 2016), HolE (Nickel et al., 2016), ConvE (Dettmers et al., 2018), RotatE (Sun et al., 2019) and MuRE (Balazevic et al., 2019). We also compare against PoLo (Liu et al., 2020) and its predecessor MINERVA (Das et al., 2018), which optimizes only the first summand of the reward function (i.e., $\lambda=0$ in Eq. 1). Additionally, we test prioritization of drug-BP triples based on degree-weighted path count (DWPC) using 0.4 as damping exponent (Himmelstein & Baranzini, 2015). We train and evaluate these models on the same data splits as MARS on MoA-net-10k.
4 Results
In an initial set of experiments on MoA-net, we observed that pruned metrics were consistently lower than standard ones (Fig. 2-A), indicating that the metapath-based rules were not being utilized in most predictions. This can happen because rule-based rewards are contingent upon a true pair being found (Eq. 1). Additionally, amongst recorded trajectories, most did not follow our metapath-based rules; instead, most trajectories used the following metapath, involving inverse edges:
This associative pattern indicates that two drugs inducing a common BP also likely induce another BP. This type of pattern is also present in a previous NeSy RL study, Liu et al. (2020), in which the most used metapath was the following:
[…]
However, when we reproduced the results from Liu et al. (2020), we achieved the same reported metrics even in the absence of the above, associative rule (See Appendix A.5). This suggests that, although the associative rule may serve as a plausible model explanation, it does not necessarily guide model training.
Previously, associative patterns have been regarded as informative in other KG prediction tasks such as DR and target prediction (Fernández-Torras et al., 2022; Ratajczak et al., 2022; Himmelstein et al., 2017; Renaux et al., 2023; Li et al., 2023; Muslu et al., 2020). However, associative patterns do not constitute biologically valid MoAs. As stated in Section 1, MoAs, like those in DrugMechDB, involve physical, molecular interactions. Therefore, an MoA path would not traverse an additional drug node. In other words, while such associations may be a valid proxy for other biomedical tasks, MoA deconvolution for novel compounds is inherently limited to information around its protein targets and related BPs.
Figure 2: Performance metrics for $\text{MARS}_{P_{2H}}$ compared to PoLo and $\text{MARS}_{\text{naive}}$ upon several variants of MoA-net. Each bar is the average and standard deviation across five independent training and testing iterations. From left to right: Little change between initial metrics upon MoA-net (A) in comparison to the standard MoA-net-permuted metrics (B) provides evidence that predictions are influenced by degree bias, resulting in a reasoning shortcut. Thereafter, inverse edges were removed to prohibit the reasoning shortcut, hindering performance (C). Performance was restored upon MoA-net-10k with the KG trimming step (D), with $\text{MARS}_{P_{2H}}$ showing the best standard and pruned metrics. Finally, $\text{MARS}_{P_{2H}}$ maintains high pruned metrics even when inverse edges (and reasoning shortcuts) are re-introduced (E).
4.2 $P_{2H}$ updates reveal reasoning shortcuts via degree bias
Previous research on KGs has shown that node degree distribution, the number of adjacent edges for each KG node, can significantly bias predictions (Tang et al., 2020; Ju et al., 2024). Specifically, inspection bias, a type of degree bias, occurs when the KG is not uniformly inspected or sampled (Zietz et al., 2024). Since PPIs are the most common relation type in MoA-net (90% of edges) (Fig. 1-B), protein nodes have a higher degree distribution than other node types. We hypothesized, therefore, that the agent circumvents denser parts of the KG, creating an inspection bias. Although rule-based predictions merit a larger reward, the MARS agent exploits associative patterns for a more reliable reward. To confirm the existence of degree bias, we tested our approach upon MoA-net-permuted. As explained in Section 3.4, MoA-net-permuted is a variant of MoA-net in which edges are swapped while preserving node degree distribution. This tests the extent to which node degree drives predictions. Indeed, the lack of change amongst standard performance metrics suggested that node degree was largely responsible for predictions (Fig. 2-B). Put simply, the agent gets lost when exploring the PPIs, so it avoids them.
4.3 Identifying and mitigating degree bias improves performance
4.5 External validation of $\text{MARS}_{P_{2H}}$ on MoA-net-10k
Table 1: Performance evaluations of MARS upon MoA-net-10k against baseline models. Metrics are presented as the (average, standard deviation) across five independent training/testing iterations. The best of each standard (top) and pruned (bottom) metric are in bold.
Model metric type Hits@1 Hits@3 Hits@10 MRR
ComplEx standard 0.137 0.291 0.527 0.263
ConvE standard 0.055 0.132 0.247 0.116
DistMult standard 0.011 0.027 0.071 0.031
ER-MLP standard 0.011 0.093 0.242 0.086
HolE standard 0.038 0.077 0.203 0.096
MuRE standard 0.071 0.181 0.253 0.177
RESCAL standard 0.055 0.148 0.286 0.135
RotatE standard 0.176 0.242 0.368 0.239
TransE standard 0.154 0.297 0.538 0.274
Metapaths with DWPC pruned 0.370 0.560 0.780 0.508
5 Discussion
NeSy approaches are sometimes portrayed as more trustworthy than their black-box counterparts, partially due to increased interpretability (Gaur & Sheth, 2024; DeLong et al., 2024). Within this study, we enhanced the interpretability of a NeSy RL approach to produce our model, $\text{MARS}_{P_{2H}}$. Specifically, through our novel algorithm, two-hop joint probabilities ($P_{2H}$), MARS learned weights corresponding to path-like rules, serving as a proxy for each rule’s importance. However, these insights revealed a new issue: NeSy RL approaches on KGs are susceptible to reasoning shortcuts. Specifically, in our study, predictions were driven by node degree bias. Ultimately, MARS’ enhanced interpretability called the trustworthiness of such approaches to question.
To address this, we considered Marconato et al. (2024a)’s desiderata for a shortcut-aware NeSy system. Specifically, on MoA-net-10k, $\text{MARS}_{P_{2H}}$ showed both competitive performance as well as calibration in comparison to other models. Notably, however, measuring calibration is challenging in this domain. While rule-based predictions, measured through pruned metrics, follow the expected semantics for MoA deconvolution, we can not determine whether every other prediction follows unintended semantics. For example, in the classic MNIST addition task, popularly used to assess NeSy methods Manhaeve et al. (2018), a model is trained to determine the sum of two handwritten digits. In this toy example, the misclassification of a handwritten ‘2’ as ‘3’ and vice versa would still amount to the same sum. Thus, reasoning shortcuts can be objectively identified. On the contrary, while we provide evidence that predictions using associative patterns are largely affected by node degree bias, we can not determine whether such patterns always reflect a reasoning shortcut.
Finally, regarding cost effectiveness, $\text{MARS}_{P_{2H}}$ can be applied to any KG, serving as a generalizeable mitigation strategy. However, we also note that this was achieved upon MoA-net-10k, a trimmed version of MoA-net. While we automated this trimming step, such a strategy does not make use of all available information. To scale $\text{MARS}_{P_{2H}}$ to denser KGs and maintain its shortcut-aware status, several future directions could be explored. For instance, one could merge similar, high-degree nodes or rely upon domain knowledge, like the identification of promiscuous proteins (Copley, 2020), to make more informed choices about edge trimming or masking.
Our study highlights a key concern in which the behavior of some NeSy RL approaches could be attributed to node degree bias, rather than meaningful, domain-specific concepts. The interpretability of our approach, $\text{MARS}_{P_{2H}}$, allowed insight into this reasoning shortcut. Therefore, we question whether such shortcuts are identifiable amongst black-box approaches. Additionally, by testing a NeSy approach upon a novel applied task, MoA deconvolution, we could flag down patterns, like associative ones, which were plausible yet arguably less meaningful to biomedical researchers. Therefore, our study emphasizes the importance of testing interpretable models, like NeSy ones, in an applied domain. Finally, while our study honors the desiderata for shortcut-aware NeSy systems, we also examined the extent to which they were applicable to a biomedical domain.
6 Conclusions
Author Contributions
LND designed the methodology, wrote the software, and conducted the investigation. YG and LND did data curation and visualization. DDF and LND conceptualized the project. DDF, JDF, and PG supervised the project. All authors took part in writing the manuscript.
Acknowledgments
LND is funded by the University of Edinburgh Informatics Graduate School through the Global Informatics Scholarship. LND partially conducted this work during an internship at Enveda Biosciences Inc. JDF and PG are funded by the National Institute for Health Research (NIHR) Artificial Intelligence and Multimorbidity: Clustering in Individuals, Space and Clinical Context (AIM-CISC) grant NIHR202639. The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care.
We thank these institutions for their support. We also thank Emile van Krieken and Guillermo Romero Moreno for their thoughtful feedback.
Appendix A Appendix
A.1 Selected Metapaths
Table A1: Metapaths representing feasible MoAs. Drugs are represented with a $D$, proteins with a $P$, and biological processes with $BP$.
$\texttt{upregulates}(D,P)\rightarrow\texttt{participates}(P,BP)$
A.2 Implementation
We implemented MARS using TensorFlow (version 2.10). The method is packaged in Python [^3] https://github.com/laurendelong21/MARS. We used a grid search hyperparameter optimization (Feurer & Hutter, 2019); further details are within Appendix A.4.
The baseline KGE models have been trained using the PyKEEN framework (v1.10.1) (Ali et al., 2021). KGEMs were trained using PyKEEN’s hyperparameter optimization pipeline over 30 trials using as initial parameters the best configurations from (Rivas-Barragan et al., 2022). The evaluation in the hyperparameter optimization was conducted using Hits@10 for all the models on a link prediction task for the previously-described splits. Network algorithms were implemented in NetworkX (v3.1) (Hagberg et al., 2008) and metapaths were calculated using the hetnetpy Python package (Himmelstein et al., 2021; Himmelstein & Baranzini, 2015). Lastly, source code and data are openly available [^4] https://github.com/laurendelong21/MoA-Net.
A.3 $P_{2H}$ formalization
Here, we describe the weight updates in more detail. Both naive and $P_{2H}$ updates are normalized according to expected values. For naive updates, we record observed frequency, $O$, at which each metapath-based rule is satisfied. This is normalized by the batch-specific expected frequency, $E$, in which every rule has a uniform probability across the total number rule groundings in that batch. For instance, if there are 10 metapath-based rules, and the agent satisfied at least one rule 100 times in a given batch, $E$ would compute to 0.10, and each observed frequency $O_{M}$, would be normalized by $E$. If the agent found zero groundings, no weight updates would be made in that batch. Ultimately, this produces a metric, $\mu$ (Eq. 3) in which any value over one indicates usefulness (the agent used that rule more than others), and any value less than one indicates otherwise.
$\mu_{M_{i}}=O_{M_{i}}/E_{M_{i}}$ (3)
Notably, the value of $\mu$ was bound to avoid division by zero and extreme values. To adjust weight updates relative to the batch size and rollouts, we defined the minimum and maximum bounds on $\mu$ as Eqs. 4 and 5, where $\rho$ is the total number of metapath-based rules:
$\mu_{min}=\frac{\rho}{\text{batch size}\times\text{rollouts}}$ (4)
$\mu_{max}=\rho\times\text{batch size}\times\text{rollouts}$ (5)
Using Eq. 6, denoted by $\Phi$, $\mu$ (Eq. 3) is used to update the weight of a rule, $w(M_{i})$. Eq. 6 is regularized by the hyperparameter $\alpha\in\mathbb{R}$, where $0\leq\alpha\leq 1$ to control how drastic the weight update is. An $\alpha$ closer to zero indicates subtle updates; this might be suitable if the user wants weights to act as a gauge on relative importance between rules. An $\alpha$ closer to one indicates drastic updates and could be suitable if the user has many rules and wants those which are less useful to be reduced to zero. If $\alpha=0$, no weight updates will be made. Therefore, $\alpha$ can be selected based on user needs or via hyperparameter optimization.
$\Phi(\mu,w(M_{i}))=w(M_{i})\times 2\alpha(\frac{\mu-1}{\mu+1})$ (6)
The second, more complex method by which weight updates were done is based on a term we coined, two-hop joint probability, $P_{2H}$. Pseudocode for $P_{2H}$ can be found in Algorithm 1 below. This metric approximates the usefulness of metapath-based rules based on the observed frequencies of the two-hop fragments which they comprise. Here, the main differences are that (1) Eq. 3 is computed based on the observed and expected probabilities of two-hop fragments, rather than whole metapaths, and (2) $\rho$ within Equations 4 and 5 is the number of unique two-hop fragments possible.
@@ A.3 $P_{2H}$ formalization
A.4 Hyperparameter selection

### v3

@@ 3.3 MARS dynamically updates rule weights
Here, the probability of each metapath-based rule is computed as the joint probability of its fragments. For example, the $P_{2H}$ metric for the above metapath-based rule would be computed as $P_{2H}(M_{example})=p(Fragment1)\times p(Fragment2)\times p(Fragment3)$. We note two caveats. Firstly, to account for partial metapath matches, we relax our definition of conjunction here, allowing truth to be evaluated on the fragment level. Secondly, metapath fragments do not necessarily represent independent events. To avoid complex computation involving conditional probabilities (Russell & Norvig, 2010), we assume independence, and the $P_{2H}$ metric serves as an approximation for the empirical probabilities of metapath-based rules.
Ultimately, MARS with $P_{2H}$ updates ($\text{MARS}_{P_{2H}}$) uses all information from successful trajectories. Compared to naive updates, $P_{2H}$ updates differ in that (1) Eq. 1 is computed based on the observed and expected probabilities of two-hop fragments, rather than whole metapaths, and (2) $\rho$ within Equations 2 and 3 is the number of unique two-hop fragments possible.
@@
Implementation details and hyperparameter selection are described in Appendices A.2 and A.6, respectively.
3.4 Datasets: MoA-net and its variants
We design our KG, MoA-net [^2] https://github.com/laurendelong21/MoA-Net, specifically for MoA prediction. MoA-net consists of drugs, proteins, and BPs (Appendix A.5). We assemble it using the causal relations between drugs and proteins from several real-world datasets comprising experimental data, including Custom KG (Rivas-Barragan et al., 2020) and OpenBioLink KG (Breit et al., 2020). The BP nodes come from experimentally-derived and expert-curated molecular function annotations in UniProt (Consortium, 2015).
To predict drug-BP triples, which are unique to MoA-net, we make use of publicly available functional and biochemical assays in ChEMBL (v33), an open access database of bioactive compounds (Gaulton et al., 2012). Of the 1,622 drug-BP triples obtained, 48 also had known MoAs in DrugMechDB (Gonzalez-Cavazos et al., 2023), a manually-curated compendium of MoAs. Between the three node types, we define five unique edge types, or relations, shown in Appendix A.5. We also include all inverse relations, running in the opposite direction of causality.
Using the hetnetpy package (Himmelstein et al., 2021), we extract all metapaths (see Section 3.1) from MoA-net which we considered to be valid MoAs: those comprising directed, mechanistic paths between drug and BP nodes (see Appendix A.1). We exclude metapaths depicting associative patterns, such as those leveraging information about shared BP targets, from our set of metapath-based rules. Based on MoAs found in DrugMechDB, we limit metapaths to a maximum length of four relations (or hops).
3.5 Evaluation
We conduct an extensive benchmark of our method against nine different baseline KG embedding (KGE) models, two state-of-the-art NeSy methods, and one network measure (Appendix A.2). We train and evaluate these models on the same data splits as MARS on MoA-net-10k.
4 Results
In an initial set of experiments on MoA-net, we observed that pruned metrics were consistently lower than standard ones (Fig. 3-A), indicating that the metapath-based rules were not being utilized in most predictions. This can happen because rule-based rewards are contingent upon a true pair being found (Eq. 5). Additionally, amongst recorded trajectories, most did not follow our metapath-based rules; instead, most trajectories used the following associative pattern, involving inverse edges:
This associative pattern indicates that two drugs inducing a common BP also likely induce another BP. This type of pattern is also present in Liu et al. (2021), in which the most used pattern was the following:
[…]
However, when we reproduced the results from Liu et al. (2021), we achieved the same reported metrics even in the absence of the above, associative rule (See Appendix A.8). This suggests that, although the associative rule may serve as a plausible model explanation, it does not necessarily guide model training. Furthermore, as stated in Section 1, MoAs, like those in DrugMechDB, involve physical, molecular interactions, rather than associative ones.
Figure 3: Hits@10 and MRR for $\text{MARS}_{P_{2H}}$ compared to PoLo and $\text{MARS}_{\text{naive}}$ upon several variants of MoA-net. Each bar is the average and standard deviation across five independent training and testing iterations. From left to right: Little change between initial metrics upon MoA-net (A) in comparison to the standard MoA-net-permuted metrics (B) provides evidence that predictions are influenced by degree bias, resulting in a reasoning shortcut. Thereafter, inverse edges were removed to prohibit the reasoning shortcut, hindering performance (C). Performance was restored upon MoA-net-10k with the KG trimming step (D), with $\text{MARS}_{P_{2H}}$ showing the best standard and pruned metrics. Finally, $\text{MARS}_{P_{2H}}$ maintains high pruned metrics even when inverse edges (and reasoning shortcuts) are re-introduced (E).
4.2 $P_{2H}$ updates reveal reasoning shortcuts via degree bias
Previous research on KGs has shown that node degree distribution, the number of adjacent edges for each KG node, can significantly bias predictions (Tang et al., 2020; Ju et al., 2024). Specifically, inspection bias, a type of degree bias, occurs when the KG is not uniformly inspected or sampled (Zietz et al., 2024). Since PPIs are the most common relation type in MoA-net (90% of edges) (Appendix A.5), protein nodes have a higher degree distribution than other node types. We hypothesized, therefore, that the agent circumvents denser parts of the KG, creating an inspection bias. Although rule-based predictions merit a larger reward, the MARS agent exploits associative patterns for a more reliable reward. To confirm the existence of degree bias, we tested our approach upon MoA-net-permuted. As explained in Section 3.4, MoA-net-permuted is a variant of MoA-net in which edges are swapped while preserving node degree distribution. This tests the extent to which node degree drives predictions. Indeed, the lack of change amongst standard performance metrics suggested that node degree was largely responsible for predictions (Fig. 3-B). Put simply, the agent gets lost when exploring the PPIs, so it avoids them.
4.3 Identifying and mitigating degree bias improves performance
4.5 External validation of $\text{MARS}_{P_{2H}}$ on MoA-net-10k
Table 1: Performance of MARS upon MoA-net-10k against baseline models. Metrics are presented as (average, standard deviation) across five independent training/testing iterations for all but DWPC, which is deterministic. The best of each standard (top) and pruned (bottom) metric are in bold.
Model metric type Hits@1 Hits@3 Hits@10 MRR
CompGCN standard (0.093, 0.010) (0.212, 0.031) (0.428, 0.043) (0.201, 0.011)
ComplEx standard (0.141, 0.025) (0.287, 0.020) (0.517, 0.007) (0.258, 0.018)
MuRE standard (0.066, 0.020) (0.157, 0.043) (0.377, 0.025) (0.160, 0.023)
PairRE standard (0.131, 0.023) (0.296, 0.035) (0.601, 0.028) (0.271, 0.022)
RotatE standard (0.132, 0.040) (0.190, 0.042) (0.349, 0.022) (0.198, 0.031)
Metapaths with DWPC pruned 0.370 0.560 0.780 0.508
5 Discussion
NeSy approaches are sometimes portrayed as more trustworthy than their black-box counterparts, partially due to increased interpretability (Gaur & Sheth, 2024; DeLong et al., 2024). Here, we presented a NeSy RL approach, $\text{MARS}_{P_{2H}}$, which promotes interpretability by deconvoluting drug MoAs. Specifically, through our novel algorithm, two-hop joint probabilities ($P_{2H}$), MARS learned weights corresponding to rules representing MoA patterns; each weight served as a proxy for each rule’s importance. However, these insights revealed a new issue: NeSy RL approaches on KGs are susceptible to reasoning shortcuts. Specifically, in our study, predictions were driven by node degree bias. Ultimately, MARS’ interpretability called the trustworthiness of such approaches to question.
To address this, we considered Marconato et al. (2024a)’s desiderata for a shortcut-aware NeSy system. Specifically, on MoA-net-10k, $\text{MARS}_{P_{2H}}$ showed both competitive performance as well as calibration in comparison to other models. Notably, however, measuring calibration is challenging in this domain. While rule-based predictions, measured through pruned metrics, follow the expected semantics for MoA deconvolution, we can not determine whether every other prediction follows unintended semantics. For example, in the classic MNIST addition task, popularly used to assess NeSy methods (Manhaeve et al., 2018), a model is trained to determine the sum of two handwritten digits. In this toy example, the misclassification of a handwritten ‘2’ as ‘3’ and vice versa would still amount to the same sum. Thus, reasoning shortcuts can be objectively identified. On the contrary, while we provide evidence that predictions using associative patterns are largely affected by node degree bias, we can not determine whether such patterns always reflect a reasoning shortcut.
Finally, regarding cost effectiveness, $\text{MARS}_{P_{2H}}$ can be applied to any KG, serving as a generalizeable mitigation strategy. However, we also note that this was achieved upon MoA-net-10k, a trimmed version of MoA-net. While we automated this trimming step, such a strategy does not make use of all available information. To scale $\text{MARS}_{P_{2H}}$ to denser KGs and maintain its shortcut-aware status, several future directions could be explored. For instance, one could merge similar, high-degree nodes or rely upon domain knowledge, like the identification of promiscuous proteins (Copley, 2020), to make more informed choices about edge trimming or masking. In addition to addressing these methodological limitations, prospective studies could explore more complex MoAs, include binding or expression values, or involve a variety of protein subclasses.
In summary, our study highlights a key concern in which the behavior of some NeSy RL approaches could be attributed to node degree bias, rather than meaningful, domain-specific concepts. The interpretability of our approach, $\text{MARS}_{P_{2H}}$, allowed insight into this reasoning shortcut. Therefore, we question whether such shortcuts are identifiable amongst black-box approaches. Additionally, by testing a NeSy approach upon a novel applied task, MoA deconvolution, we could flag down patterns, like associative ones, which were plausible yet arguably less meaningful to biomedical researchers. Therefore, our study emphasizes the importance of testing interpretable models, like NeSy ones, in an applied domain. Finally, while our study honors the desiderata for shortcut-aware NeSy systems, we also examined the extent to which they were applicable to a biomedical domain.
6 Conclusions
Appendix A Appendix
A.1 Selected Metapaths
Table A1: Metapaths representing MoAs. Drugs are represented with a $D$, proteins with a $P$, and biological processes with $BP$.
$\texttt{upregulates}(D,P)\rightarrow\texttt{participates}(P,BP)$
We use the MoAs in DrugMechDB (Gonzalez-Cavazos et al., 2023) as guidance for the types of MoA patterns which should exist within our selected metapaths. To get MoAs most relevant for our study, we extracted paths between drugs and BPs within DrugMechDB. All of such paths were $\leq 4$ hops long, justifying the maximum length of paths in Table A1:
Figure 5: DrugMechDB paths extracted between drugs and BPs.
A.2 Implementation
We implemented MARS using TensorFlow (version 2.10). The method is packaged in Python [^3] https://github.com/laurendelong21/MARS. The neural network structure is implemented as in Liu et al. (2021), which is also drawn from Das et al. (2018). We used the Adam optimizer (Kingma & Ba, 2015) with REINFORCE (Williams, 1992) to maximize rewards. We used a grid search hyperparameter optimization (Feurer & Hutter, 2019); further details are within Appendix A.6. MARS was trained to optimize MRR, with early stopping determined by validation MRR (Fig. 6).
Figure 6: Validation MRR over training epochs.
Partially based on previous biomedical KG benchmarks (Rivas-Barragan et al., 2022), the KGE baseline models include ComplEx (Trouillon et al., 2016), RotatE (Sun et al., 2019), MuRE (Balazevic et al., 2019), CompGCN (Vashishth et al., 2020), and PairRE (Chao et al., 2021). We also compare against PoLo (Liu et al., 2021) and its predecessor MINERVA (Das et al., 2018), which is not guided by rules. Additionally, we test prioritization of drug-BP triples based on degree-weighted path count (DWPC) using 0.4 as damping exponent (Himmelstein & Baranzini, 2015).
The baseline KGE models were trained using the PyKEEN framework (v1.10.1) (Ali et al., 2021). KGEMs were trained using PyKEEN’s hyperparameter optimization pipeline over 30 trials using as initial parameters the best configurations from (Rivas-Barragan et al., 2022). The evaluation in the hyperparameter optimization was conducted using Hits@10 for all the models on a link prediction task for the previously-described splits. Network algorithms were implemented in NetworkX (v3.1) (Hagberg et al., 2008) and metapaths were calculated using the hetnetpy Python package (Himmelstein et al., 2021; Himmelstein & Baranzini, 2015). Lastly, source code and data are openly available [^4] https://github.com/laurendelong21/MoA-Net.
A.3 Hardware and Resources
For training MARS, we used one A40 Nvidia GPU (NVIDIA Corporation, 2021) and two AMD EPYC Milan 7413 CPU nodes (AMD, 2021). On MoA-net, MARS used up to 90 Gigabytes of memory and took up to 1.5 hours between beginning training and concluding testing.
A.4 Reward function
Using a deep RL process, MARS trains an agent to take walks of length $L$ through the KG to connect pairs of nodes having the pre-defined relation of interest (e.g., induces); such edges are masked from the agent during training. Each walk generates a path, $P$, which can be understood as a series of $L$ transitions: $P:=(e_{c}\xrightarrow{r_{1}}e_{2}\xrightarrow{r_{2}}...\xrightarrow{r_{L}}e_{L+1})$. The agent may also remain at its current node. Ultimately, the goal of the agent is episodic: to find paths in which the starting node, $e_{c}$, and the terminal node, $e_{L+1}$, make up one of the true input pairs. By training the agent to do so, it can identify node pairs with some desired relationship, thus generalizing beyond the training set to predict novel pairs. Through a process akin to a Markov Decision Process (Bellman, 1957), the agent makes decisions about its next move based on information about its current position and the next possible actions. Additionally, however, the history of the agent’s previous actions are encoded with an LSTM (Hochreiter & Schmidhuber, 1997; Sherstinsky, 2020), whose parameters are trained to optimize the reward function, $R(S_{L+1})$ (Eq. 5), which is evaluated each time the agent completes $L$ transitions from some starting node, reaching a state $S_{L+1}$. MARS uses the same reward function as Liu et al. (2021):
This reward function which quantifies how successful $P$ is according to two summands, where the hyperparameter $\lambda$ influences the balance between them. The first summand indicates whether the terminal node in the path, $e_{L+1}$, is one of the desired target (BP) nodes ($e_{d}$) that forms a true pair with the starting (drug) node, $e_{c}$. The second summand, contingent upon the first, indicates whether the corresponding metapath, $\tilde{P}$, matches any metapath-based rule, $M_{i}$. Therefore, the agent is not only encouraged to find connections between true pairs of nodes, but it is also guided toward paths which resemble known MoAs. Of note, the second summand is proportional to some metapath-based rule weight, which is learned by MARS.
A.5 Node and edge type distribution
Node type Count
$Drug$ 300
$Protein$ 9,301
$Biological\ Process\ (BP)$ 86
Edge type Count
$\texttt{interacts}(Protein,Protein)$ 86,786
$\texttt{participates}(Protein,BP)$ 4,325
$\texttt{downregulates}(Drug,Protein)$ 2,205
$\texttt{upregulates}(Drug,Protein)$ 1,631
$\texttt{induces}(Drug,BP)$ 1,622
A.6 Hyperparameter selection